# Chip-Scale Point-Source Sagnac Interferometer by Phase-Space Squeezing

### v2

@@ Discussion
To examine the impact of the SPSI, we conduct detailed numerical simulations for PSI and SPSI. Initially, we generate a cloud of $N$ atoms following the cooling phase, characterized by a spatial Gaussian distribution with a standard deviation of $\sigma_{x0}$. The atoms’ velocities follow a Maxwell-Boltzmann distribution corresponding to the given temperature. In a standard PSI simulation, the cloud undergoes expansion, followed by the initiation of the atomic-interferometer sequence with a time interval $T_{R}$ between pulses. Upon completion of this sequence, we determine the internal state of each atom using a Monte Carlo method. We simulate a two-state absorption of the atomic cloud by counting the atoms of each state in bins of spatial locations and normalizing the counts of one state to the total population in each bin $P_{1}=N_{|1\rangle}(x)/N_{\rm total}(x)$. This approach reveals the interferometer’s fringe pattern as per Eq. 2, illustrated in Fig. 5. Utilizing a sinusoidal fit allows us to extract the fringe’s periodicity, $k_{x}$, facilitating the calculation of angular velocity based on Eq. 3. The $\alpha$ factor is accounted for by scaling the fit’s frequency accordingly before extracting the angular velocity as described in Eq. 2. The uncertainty in angular velocity, denoted as single-shot sensitivity $\delta\Omega$, is computed based on the fit uncertainty. Running simulations for multiple repetitions and calculating the standard deviation provides a comparable result to the uncertainty derived from a single run’s fit.
In an SPSI simulation, we incorporate an additional stage: accelerating the atoms with a repulsive potential. We employ a blue-detuned inverted harmonic shaped light potential (Eq. 11) with a total power $P$ and beam dimensions $x_{0}$ and $z_{0}$, which accelerates the atoms over a brief duration of $t_{\rm acc}$. Following the acceleration stage, the simulation resumes in a manner identical to a standard PSI.
The normalized population is fitted to a sinusoidal function, $f(x)=A\sin(Bx+C)+D$, which initially employs an FFT algorithm to identify optimal frequencies and uses them as initial guesses for the fit algorithm, based on non-linear least-squares. To improve the fit’s success rate, we implement a basic image processing stage before normalizing $P_{1}$. This stage involves applying different averaging windows to the image’s pixels to reduce noise in the atom count in each bin. Additionally, the algorithm examines various windows of interest from the image, excluding the noisier edges where fewer atoms are counted and more fluctuations occur.
@@ Acknowledgments
The preparation sequence for the atomic cloud begins with a 3D magneto-optical trap, followed by optical molasses to reduce the cloud’s temperature. Prior to applying the repulsive potential, we optically pump the 87Rb atoms to the $|F=2,m_{F}=2\rangle$ atomic state. The repulsive potential is blue-detuned with circularly polarized $\sigma^{+}$ light, which ensures that any undesired excitation of the atoms decays back to the $|F=2,m_{F}=2\rangle$ state. The repulsive potential defines the quantization axis, eliminating the need for external magnetic fields. At the end of the acceleration and just before the interferometer sequence, we pump the atoms to the non-magnetic $|F=1,m_{F}=0\rangle$ state, so that the interferometric sequence may begin.
The blue-detuned light induces a positive-energy shift of the ground atomic level and a negative-energy shift of the excited level. As the energy shift is proportional to the light intensity, an atom in the ground level is repelled from the high-intensity beam center, while the excited level is drawn towards it. For the repulsion to be effective and to avoid heating by spontaneous emission, it is essential that $\Omega_{R}^{2}\ll\Delta^{2}$, so that the excited-level occupation stays small. For example, for the repulsive potential parameters used in Fig. 2, in the center of the potential ($I(x,z)=I_{0}$) we get $\Omega_{R}=2\pi\cdot 3.16\,\rm{GHz}$, such that $\Delta=2\pi\cdot 10\,\rm{GHz}$ gives a ratio of $\Delta^{2}/\Omega_{R}^{2}=10$.
@@
The repulsive potential force in the direction of optical beam propagation adds a velocity in this axis equal to $v_{\text{long}}=N_{\text{sc}}v_{\text{rec}}$. Using the same potential and interferometer parameters as in Fig. 2, $v_{\text{long}}$ is equal to $28\,\rm{cm/s}$, resulting in a total longitudinal displacement of $\Delta x_{\text{long}}=2.8\,\rm{mm}$. This can be easily addressed by enlarging the vacuum cell or by setting the initial cloud away from the center of the cell.
@@
In a two-dimensional phase space, of position $x$ and velocity $v$, a Gaussian distribution in the form of an exponent of a bi-quadratic expression in $x$ and $v$ remains Gaussian in either free evolution or evolution in a quadratic potential. This conservation of the Gaussian form follows from the fact that any linear transformation of a bi-quadratic expression of two variables remains bi-quadratic under such a transformation. Any Gaussian form where the distribution is centered around $x=0$ and $v=0$ can be written in the form
$\rho(x,v,t)\propto e^{-(x-vt)^{2}\sigma_{x0}^{2}}e^{-v^{2}/2\sigma_{v0}^{2}}\,,$ (15)
@@
Consider an initial wave-function $\psi({\bf r})$ subject to a sequence of splitting and recombining pulses $\pi/2-\pi-\pi/2$ with an initial state $|0\rangle$ and another state $|1\rangle$. In the inertial frame, the initial pulse applies a momentum transfer $K_{\text{eff}}$ in the $\hat{z}$ direction to the state $|1\rangle$. The second pulse applies momentum kicks ±$\pm K_{\text{eff}}$ in a rotated direction $\cos\Omega T_{R}\hat{z}+\sin\Omega T_{R}\hat{x}$ to the two states, and the third pulse applies the positive momentum to the $|1\rangle$ state in the rotated direction $\cos 2\Omega T_{R}\hat{z}+\sin 2\Omega T_{R}\hat{x}$. When $\Omega T_{R}\ll 1$, we can ignore the effect of the rotation in the $\hat{z}$ direction, as $\cos\Omega T_{R}\approx 1$, and consider only the dynamics along the $\hat{x}$ direction, where we can approximate $\sin\Omega T_{R}\approx\Omega T_{R}$. Then the two paths end up in the same velocity but in different positions: The $|1\rangle-|0\rangle$ pulse ends up in $x=x_{0}+2v_{0}T_{R}-v_{R}\Omega T_{R}^{2}$ and the $|0\rangle-|1\rangle$ paths ends up in $x_{0}+2v_{0}T_{R}+v_{R}\Omega T_{R}^{2}$, where $v_{R}=\hbar K_{\text{eff}}/m$ is the recoil velocity. The phase difference between the two paths due to the velocity difference in the $x$-direction follows from the rotation
$\delta\phi_{\rm rot}=\frac{m}{2\hbar}[(v_{0}-v_{R}\Omega T_{R})^{2}-(v_{0}+v_{R}\Omega T_{R})^{2}]T_{R}=-2K_{\text{eff}}v_{0}\Omega T_{R}^{2}.$ (20)
@@
The wave function in the state $|0\rangle$ after the sequence is then
$\psi_{f}(x)=\frac{1}{\sqrt{2}}[\psi_{\sigma}(x-x_{a})+\psi_{\sigma}(x-x_{b})e^{-i\delta\phi_{\rm rot}}],$ (21)
where $x_{a}$ and $x_{b}$ are the final positions of the two paths and $\psi_{\sigma}$ is the wave function after expansion from an initial width $\sigma_{0}$ to a final width $\sigma$ after a time $t=2T_{R}$. It follows that the final wave function is similar to the result of an expansion starting with two Gaussian wave-packets centered at a relative distance $d=|x_{a}-x_{b}|=2v_{R}\Omega T_{R}^{2}$, leading to the emergence of a fringe pattern in a manner equivalent to the double-slit experiment. Here, the distance $d$ is proportional to the square of the interferometer time, but the fringe periodicity depends on the total expansion time since the wave packet had a minimal size.
@@
The evolution of a Gaussian wave-packet in free space, or alternatively a Bose-Einstein wave-packet in free space or a quadratic potential, can be expressed in terms of the evolution of the width $\sigma$. The simplified expression for one dimension in the frame moving with the center of the wave-packet is given by  (?, ?)
@@
where $\psi_{0}(x)$ is the initial wave-packet wave function and $\lambda(t)=\sigma(t)/\sigma(0)$. The evolution of the width of a Gaussian wave-packet in free space is given by
$\sigma(t)=\sqrt{\sigma(0)^{2}+\hbar^{2}t^{2}/4m^{2}\sigma(0)^{2}}=\sqrt{\sigma_{x0}^{2}+\sigma_{v0}^{2}t^{2}}\,,$ (23)
with $\sigma_{v0}=\hbar/2m\sigma_{x0}$ for a coherent Gaussian wave-packet.
If we take, without loss of generality, $x_{0}=0$ and $v_{0}=0$ then the phase $\delta\phi_{\rm rot}=0$ and the phase difference between the two wave functions centered at $x_{a}$ and $x_{b}$ at a given point $x$ follows from the quadratic phases of the expanding Gaussians of the form of Eq. 22 and given by
$\delta\phi(x)=\frac{m\dot{\sigma}}{2\hbar\sigma}[(x-v_{R}\Omega T_{R}^{2})^{2}-(x+v_{R}\Omega T_{R}^{2})^{2}]=-k_{\Omega}x,$ (24)
where $k_{\Omega}=2(\dot{\sigma}/\sigma)K_{\text{eff}}\Omega T_{R}^{2}$.
For a Gaussian wave-packet with an initial width $\sigma_{0}$, the relative rate of expansion is given by
$\frac{\dot{\sigma}}{\sigma}=\frac{\omega^{2}t}{1+\omega^{2}t^{2}}=\frac{1}{t}\left(1-\frac{\sigma_{x0}^{2}}{\sigma_{f}^{2}}\right)\,,$ (25)
where $\omega=\hbar/2m\sigma(0)^{2}=\sigma_{v0}/\sigma_{x0}$. Therefore, we obtain the semi-classical result for expansion time of $t=T_{ex}$
$k_{\Omega}=2K_{\text{eff}}\Omega\frac{T_{R}^{2}}{T_{ex}}\left(1-\frac{\sigma_{0}^{2}}{\sigma_{f}^{2}}\right).$ (26)
In the limit of long times, where $\sigma_{f}\gg\sigma_{0}$, this becomes the well-known double-slit interference wave vector $k_{\Omega}=md/\hbar T_{ex}$. The contrast is obtained by observing the amplitude of the interference term
@@
$C=e^{-d^{2}/8\sigma_{f}^{2}}=\exp\left(-\frac{1}{2}k_{\Omega}^{2}\frac{\sigma_{0}^{2}}{1-\sigma_{0}^{2}/\sigma_{f}^{2}}\right).$ (28)
@@
where $n$ is the number of nodes in the function and runs from 00 to $\infty$ and $w_{n}\propto\exp(-n\hbar\omega/k_{B}T)$ are the weights (probabilities) of the states $|n\rangle$. It is clear that a thermal Boltzmann distribution yields a Gaussian spatial distribution $\sum_{n}w_{n}|\Phi_{n}(x)|^{2}\propto\exp(-x^{2}/2\sigma_{T}^{2})$ with $\sigma_{T}=\sqrt{k_{B}T/m\omega^{2}}$ in a harmonic trap. It follows that for a thermal distribution $\sigma_{0}$ in Eq. 28 can be replaced by the width of the thermal distribution to obtain the right expression for the contrast of an interferometer with a thermal cloud. This expression for the contrast is identical to the expression appearing in  (?).

### v3

@@ Discussion
To examine the impact of the SPSI, we conduct detailed numerical simulations for PSI and SPSI. Initially, we generate a cloud of $N$ atoms following the cooling phase, characterized by a spatial Gaussian distribution with a standard deviation of $\sigma_{x0}$. As we deal exclusively with thermal clouds, the atoms’ velocities follow a Maxwell-Boltzmann distribution corresponding to the given temperature. In a standard PSI simulation, the cloud undergoes expansion, followed by the initiation of the atomic-interferometer sequence with a time interval $T_{R}$ between pulses. Upon completion of this sequence, we determine each atom’s internal state via Eq. 2, utilizing a Monte Carlo method. The wave-vector $k_{v}$ and the phase $\phi_{s}$ are determined by the rotation angular velocity (parameter $\Omega$) as in Eq. 2. We simulate absorption imaging of the two atomic states across the cloud by counting the atoms of each state in bins of spatial locations and normalizing the counts of one state to the total population in each bin $P_{1}(x)=N_{|1\rangle}(x)/N_{\rm total}(x)$. This approach reveals the interferometer’s fringe pattern as per Eq. 3, illustrated in Fig. 5. Utilizing a sinusoidal fit allows us to extract the fringe’s periodicity, $k_{x}$, facilitating the calculation of angular velocity based on Eq. 4. The $\alpha$ factor is accounted for by scaling the fit’s frequency accordingly before extracting the angular velocity as described in Eq. 4. The uncertainty in angular velocity, denoted as single-shot sensitivity $\delta\Omega_{\text{single}}$, is computed based on the fit uncertainty. Running simulations for multiple repetitions and calculating the standard deviation provides a comparable result to the uncertainty derived from a single run’s fit.
In an SPSI simulation, we incorporate an additional stage: accelerating the atomic ensemble via a repulsive potential. We implement a blue-detuned inverted harmonic-shaped optical potential (Eq. 13) characterized by a total power $P$ and beam dimensions $\Delta x_{rep}$ and $\Delta z_{rep}$, applied over a brief duration of $t_{\rm acc}$. By numerically solving the atomic dynamical system using an ordinary differential equation (ODE) solver, we quantify the acceleration’s kinematic effects. Upon completion of the acceleration stage, the simulation protocol proceeds consistently with a standard PSI methodology.
The normalized population is fitted to a sinusoidal function, $f(x)=A\sin(Bx+C)+D$, which initially employs an FFT algorithm to identify optimal frequencies and uses them as initial guesses for the fit algorithm based on non-linear least-squares. To improve the fit’s success rate, we implement a basic image processing stage before normalizing $P_{1}$. This stage involves applying different averaging windows to the image’s pixels to reduce noise in the atom count in each bin. Additionally, the algorithm examines various windows of interest from the image, excluding the noisier edges where fewer atoms are counted and more fluctuations occur.
@@ Acknowledgments
In this section, we provide more details about the preparation of the atoms and the repulsive potential and discuss the effects of heating induced by the light beam responsible for the repulsive potential. The preparation sequence for the atomic cloud begins with a 3D magneto-optical trap, followed by optical molasses to reduce the cloud’s temperature. Before applying the repulsive potential, the 87Rb atoms are optically pumped to the $|F=2,m_{F}=2\rangle$ atomic state. The repulsive potential is blue-detuned with circularly polarized $\sigma^{+}$ light, which ensures that any undesired excitation of the atoms decays back to the $|F=2,m_{F}=2\rangle$ state. The repulsive potential defines the quantization axis, eliminating the need for external magnetic fields. At the end of the acceleration and just before the interferometer sequence, the atoms are pumped to the non-magnetic $|F=1,m_{F}=0\rangle$ state so that the interferometric sequence may begin.
The blue-detuned light induces a positive-energy shift of the ground atomic level and a negative-energy shift of the excited level. As the energy shift is proportional to the light intensity, an atom in the ground level is repelled from the high-intensity beam centre while the excited level is drawn towards it. For the repulsion to be effective and to avoid heating by spontaneous emission, it is essential that $\Omega_{R}^{2}\ll\Delta^{2}$ so that the excited-level occupation stays small. For example, for the repulsive potential parameters used in Fig. 2, in the centre of the potential ($I(x,z)=I_{0}$) we get $\Omega_{R}=2\pi\cdot 3.16\,\rm{GHz}$, such that $\Delta=2\pi\cdot 10\,\rm{GHz}$ gives a ratio of $\Delta^{2}/\Omega_{R}^{2}=10$.
@@
The repulsive potential force in the direction of optical beam propagation adds a velocity in this axis equal to $v_{\text{long}}=N_{\text{sc}}v_{\text{rec}}$. Using the same potential and interferometer parameters as in Fig. 2, $v_{\text{long}}$ is equal to $28\,\rm{cm/s}$, resulting in a total longitudinal displacement of $\Delta x_{\text{long}}=2.8\,\rm{mm}$. This can be easily addressed by enlarging the vacuum cell or by setting the initial cloud away from the centre of the cell.
@@
In this section we describe the evolution of a Gaussian position-momentum phase-space distribution due to free propagation, which is valid for both classical and quantum systems, and discuss the effect of heating on such a distribution. In a two-dimensional phase space of position $x$ and velocity $v$, a Gaussian distribution in the form of an exponent of a bi-quadratic expression in $x$ and $v$ remains Gaussian in either free evolution or evolution in a quadratic potential. This conservation of the Gaussian form follows from the fact that any linear transformation of a bi-quadratic expression of two variables remains bi-quadratic under such a transformation. Any Gaussian form where the distribution is centred around $x=0$ and $v=0$ can be written in the form
$\rho(x,v,t)\propto e^{-(x-vt)^{2}/2\sigma_{x0}^{2}}e^{-v^{2}/2\sigma_{v0}^{2}}\,,$ (17)
@@
Although this work is concerned with thermal atomic clouds with classical evolution, we add here a description of the PSI interferometer from a quantum perspective. The description is based on the wave packet approach of Ref. (?) and may be extended back to the classical regime by treating a thermal cloud of atoms with a density matrix diagonal in the basis of hermite-Gaussian states. This approach was not used in the main text of the paper but may provide a starting point for extension to the quantum regime if a BEC is used as the atomic source.
@@
Consider an initial wave-function $|\psi(t=0)\rangle=\psi({\bf r},t=0)|0\rangle$ in an initial internal state $|0\rangle$ subject to a sequence of Raman pulses $\pi/2-\pi-\pi/2$ with one arm of the interferometer going through the state sequence $|0\rangle\to|1\rangle\to|0\rangle\to|1\rangle$ and the other through $|0\rangle\to|0\rangle\to|1\rangle\to|1\rangle$. In the inertial frame the initial $\pi/2$ pulse applies a momentum transfer $k_{\rm eff}$ in the $\hat{z}$ direction to the first arm. The second pulse ($\pi$ pulse) applies momentum kicks $\mp k_{\rm eff}$ in a rotated direction $\cos\Omega T_{R}\hat{z}+\sin\Omega T_{R}\hat{x}$ to the two arms, and the last $\pi/2$ pulse applies the positive momentum to the first arm in the rotated direction $\cos 2\Omega T_{R}\hat{z}+\sin 2\Omega T_{R}\hat{x}$. When $\Omega T_{R}\ll 1$ we can ignore the effect of the rotation in the $\hat{z}$ direction, as $\cos\Omega T_{R}\approx 1$, and consider only the dynamics along the $\hat{x}$ direction, where we can approximate $\sin\Omega T_{R}\approx\Omega T_{R}$. The two classical paths representing the wave-packet centers end up in the same momentum $p_{x}=\hbar k_{\rm eff}\Omega T_{R}$ but in different positions: The first and second arms end up at $x_{\mp}=x_{0}+2v_{0}t_{R}\mp v_{R}\Omega T_{R}^{2}$, where $v_{R}=\hbar k_{\rm eff}/m$ is the recoil velocity. The difference between the phases at the wave-packet centers due to the velocity difference in the $x$-direction follows from the rotation
$\delta\phi_{\rm rotation}=\frac{m}{2\hbar}\left[(v_{0}-v_{R}\Omega T_{R})^{2}-(v_{0}+v_{R}\Omega T_{R})^{2}\right]T_{R}=-2k_{\rm eff}v_{0}\Omega T_{R}^{2}\,.$ (22)
@@
The wave function in the state $|1\rangle$ after the sequence (time $t_{f}$) is then
$\langle 1|\psi(t_{f})\rangle=\frac{1}{\sqrt{2}}[\psi_{\sigma}(x-x_{-},t_{f})+\psi_{\sigma}(x-x_{+},t_{f})e^{-i\delta\phi}],$ (23)
where $\delta\phi$ is the phase difference between the two arms, including rotation and gravity contributions, and $\psi_{\sigma}(x,t)$ is the wavefunction (reduced to one dimension for simplicity) after expansion from an initial width $\sigma_{0}$ to a final width $\sigma(t)$ after a time $t=t_{f}=T_{\rm ex}$.
It follows that the final wave function is similar to the result of an expansion starting with two wave-packets centered at a relative distance
$d=|x_{-}-x_{+}|=2v_{R}\Omega T_{R}^{2}\,.$ (24)
After the expansion this leads to the emergence of a fringe pattern in a manner equivalent to the double-slit experiment. Here the distance $d$ is proportional to the square of the interferometer time and the fringe periodicity in the limit of long expansion time is then given by the far-field limit of the double-slit interference,
$k_{\Omega}^{PS}=\frac{md}{\hbar T_{\rm ex}}=\frac{2k_{\rm eff}\Omega T_{R}^{2}}{T_{\rm ex}}\,,$ (25)
which is identical to the point-source limit of Eq. (3) of the main text.
@@
After deriving the point-source limit, we are going to obtain the full details of the signal for a finite-size wave packet expanding for a finite time. The evolution of a Gaussian wave-packet in free space, or alternatively a BEC wave-packet in free space or, more generally, in a quadratic potential can be expressed in terms of the evolution of the width $\sigma(t)$. The expression for one dimensional free-space propagation in the frame moving with the center of the wave-packet is given by (?, ?)
@@
where $\psi_{0}(x)$ is the initial wave-packet wave function and $\lambda(t)=\sigma(t)/\sigma(0)$ is the scaling factor of the wave-packet width. The evolution of the width of a Gaussian wave-packet in free space is given by
$\sigma(t)=\sqrt{\sigma(0)^{2}+\frac{\hbar^{2}t^{2}}{4m^{2}\sigma(0)^{2}}}=\sqrt{\sigma_{x0}^{2}+\sigma_{v0}^{2}t^{2}}\,,$ (27)
with $\sigma_{x0}\equiv\sigma(0)$ and $\sigma_{v0}=\hbar/2m\sigma_{x0}$ for a coherent Gaussian wave-packet, whose momentum uncertainty is proportional to the inverse of the position uncertainty $\sigma(0)$.
If we take, without loss of generality, $x_{0}=0$ and $v_{0}=0$ then the rotational phase $\delta\phi_{\rm rotation}$ of Eq. (22) vanishes. Here we also ignore the gravitational phase due to the propagation along $z$. The phase difference between the two wave functions centered at $x_{\mp}$ at a given point $x$ follows from the quadratic phases of the expanding Gaussians of the form of Eq. (26) and given by
$\delta\phi(x)=\frac{m\dot{\sigma}}{2\hbar\sigma}[(x-d/2)^{2}-(x+d/2)^{2}]=-k_{\Omega}x,$ (28)
where
$k_{\Omega}=\frac{m}{\hbar}\frac{\dot{\sigma}}{\sigma}d=k_{\omega}^{PS}\alpha\,,$ (29)
where $\alpha=\dot{\sigma}T_{\rm ex}/\sigma$ with $\sigma$ given at $t=T_{\rm ex}$. By taking $\sigma(t)$ to be given in Eq. (27) we obtain
$\alpha=\frac{\sigma_{v0}^{2}T_{\rm ex}^{2}}{\sigma_{x0}^{2}+\sigma_{v0}^{2}T_{\rm ex}^{2}}=1-\frac{\sigma_{x0}^{2}}{\sigma_{f}^{2}}\,,$ (30)
which is identical to the definition of $\alpha$ after Eq. (4) of the main text.
The contrast is obtained from the amplitude of the interference term
@@
$C=e^{-d^{2}/8\sigma_{f}^{2}}=\exp\left[-\frac{1}{2}\alpha(k_{\Omega}^{PS}\sigma_{x0})^{2}\right]\,,$ (32)
where we have used $d=(\hbar/m)k_{\Omega}^{PS}T_{\rm ex}=2\sigma_{x0}\sigma_{v0}T_{\rm ex}k_{\Omega}^{PS}$ from Eq. (25) and the minimum uncertainty relation $m\sigma_{v0}\sigma_{x0}=\hbar/2$. Eq. (32) is identical with Eq. (5) of the main text and we have therefore completed the proof of the identity of the quantum signal with the semi-classical results.
@@
where $n$ is the number of nodes in the function and runs from 00 to $\infty$ and $w_{n}\propto\exp(-n\hbar\omega/k_{B}T)$ are the weights (probabilities) of the states $|n\rangle$. The wave function of each of these states is a product of a Gaussian with the same width $\sigma$ for all $n$ and a Hermite polynomial and the evolution of these wave functions in free space follows the same scaling as the state $|n=0\rangle$ analyzed above. A thermal Bolzmann distribution yields a Gaussian spatial distribution $\sum_{n}w_{n}|\Phi_{n}(x)|^{2}\propto\exp(-x^{2}/2\sigma_{T}^{2})$ with $\sigma_{T}=\sqrt{k_{B}T/m\omega^{2}}$ in a harmonic trap. It can be shown that the signal resulting from an interferometric sequence that uses a thermal cloud is then identical to the signal resulting from the coherent wave packet, with the position and momentum widths $\sigma_{x0}$ and $\sigma_{v0}$ replaced by the thermal phase-space widths. These results agree completely with the results that follow from the semi-classical derivation that was used in the main text.